The cooling system in a truck contains 4 gallons of coolant that is 30% antifreeze. How much must be drained and replaced with 100% antifreeze to bring the coolant in the system to 50% antifreeze?
Answer by mananth(16946)

let 100 % required be x gallons
30 % 4 gallons
Mixture 50% x+4 gallons
...
x+0.3*4= 0.5(x+4)
x+1.2=0.5x+2
x-0.5x=2-1.2
0.5x=0.8
x= 0.8/0.5
=1.6 gallons has to be drained and replaced